OpenStudy (anonymous):

Question on displacement & velocity. Will post the question and my idea on how to do it.

OpenStudy (anonymous):

on the first one "[3,4]" I tried the answer 4. from: s = t2 − 6t + 18 s' = 2t - 6 s' = 2(3) - 6 m = 0 y - 4 = (0)(x-x1) y = 4 but it didnt work. do i just substitute the values 3 and 4 on the original equation then add them then divide by 2 to get the average?

OpenStudy (anonymous):

wiat scratch that. so... do i use s' = 2t - 6 and substitute the values like: s' = 2(3) - 6 s' = 0 m/s and s' = 2(4) - 6 s' = 8-6 s' = 2 and average the two?

OpenStudy (anonymous):

\[\frac{ s(4) -s(3) }{ 4-3}\]

OpenStudy (anonymous):

really? that's it?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

k ill try it

OpenStudy (anonymous):

your second method works too... but I get the feeling they want you to do this the way I did it...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

haha! thanks! :D

OpenStudy (anonymous):

sure:)